# Large Language Model Confidence Estimation via Black-Box Access

### v2

@@ Abstract
Estimating uncertainty or confidence in the responses of a model can be significant in evaluating trust not only in the responses, but also in the model as a whole. In this paper, we explore the problem of estimating confidence for responses of large language models (LLMs) with simply black-box or query access to them. We propose a simple and extensible framework where, we engineer novel features and train a (interpretable) model (viz. logistic regression) on these features to estimate the confidence. We empirically demonstrate that our simple framework is effective in estimating confidence of Flan-ul2, Llama-13b and Mistral-7b on four benchmark Q&A tasks as well as of Pegasus-large and BART-large on two benchmark summarization tasks with it surpassing baselines by even over $10\%$ (on AUROC) in some cases. Additionally, our interpretable approach provides insight into features that are predictive of confidence, leading to the interesting and useful discovery that our confidence models built for one LLM generalize zero-shot across others on a given dataset.
@@ 1 Introduction
Having black-box access to an LLM limits the strategies one could leverage to ascertain confidence, but if the proposed strategies are effective they could be widely applied. Previous approaches [16, 22, 12] predominantly exploit the variability in the outputs for a given input prompt or based on an ensemble of prompts computing different estimators. Our approach enhances this idea where we design different ways of manipulating the input prompt and based on the variability of the answers produce values for each such manipulation. We aver to these values as a features. Based on these features computed for different inputs we train a model to predict if the response was correct or incorrect. The probability of each such prediction is then the confidence that we output. Since, the models we use to produce such predictions are simple (viz. logistic regression) the confidence estimates are typically well calibrated [24]. Moreover, being interpretable we can also see which features were more crucial in the estimation. This general framework and the features we engineer are shown in Figure 1. The framework is extensible, since more features or prompt perturbations can be easily added to this framework.
@@ 2 Related Work
Table 1: Below we see examples of different prompt perturbations for a prompt from the SQuAD dataset. The color blue and strike outs indicate changes to the input prompt. i) SD does not change the prompt (hence empty cell), but using a stochastic decoding scheme samples multiple responses (four example samplings shown). PP paraphrases the prompt. SP randomly reorders some of the sentences. EFA repeats certain sentences with entities in them. SR removes stopwords. SRC checks for consistency in reasonable size random splits of the LLM response (again prompt is not perturbed). The splitting of the two sentences indicates inconsistency as depicted in red. Thus, the perturbations test an LLM in complimentary ways.
@@ 3.1 Prompt Perturbations
Paraphrasing (PP): In this strategy we paraphrase the context in the prompt and observe how that changes the output. An example of this is shown in Table 1. For paraphrasing, we use back translation, where we convert the original prompt into another language and translate it back into English. We use machine translation models from Helsinki-NLP on huggingface and translate the text from English to French and then back to English. This new prompt then can be used to query the LLM. Changes to the output could indicate brittleness in the LLMs original response. One could also prompt an LLM to paraphrase the responses, however, in our initial experiments, we observed that when context is involved, the model does not paraphrase the entire context and parts of it are omitted.
@@ 3.2 Featurization
Semantic Set: Based on the responses of the first strategies (run multiple times) we create semantically equivalent sets for each. A semantically equivalent set consists of outputs that are semantically the same. If an response entails another response and vice-versa, then they both are grouped under the same semantic set. The number of such sets is a feature for our model. As such, more the number of sets lower the confidence estimate.
Lexical Similarity: We compute the average lexical similarity for outputs of each of the first five strategies (run multiple times). The similarity can be measured using standard NLP metrics such as rouge, blue score etc. The higher the lexical similarity higher the estimated confidence. We use rouge score to quantify the lexical similarity.
SRC Minimum Value: As mentioned above, semantic inconsistency between the two parts is measured using an NLI models contradiction probability, where one part is taken as the premise and the other the hypothesis. The highest contradiction probability amongst multiple such samplings is the feature value for this strategy.
@@ 3.3 Label Creation and Confidence Estimation
Table 2: AUROCs on four Q&A and two summarization datasets (CNN, XSUM) using a total of five LLMs (Llama, Flan-ul2, Mistral, Pegasus, BART). Higher values are better. Best results bolded.
Dataset(LLM) # of Semantic Lexical EigenValue Eccentricity Degree Ours
Sets Similarity
TriviaQA(Llama) 0.73 0.76 0.77 0.76 0.77 0.88
TriviaQA(Flan-ul2) 0.83 0.8 0.86 0.86 0.87 0.95
TriviaQA(Mistral) 0.65 0.72 0.76 0.75 0.75 0.81
SQuAD(Llama) 0.65 0.72 0.74 0.58 0.72 0.83
SQuAD(Flan-ul2) 0.6 0.7 0.67 0.65 0.67 0.8
SQuAD(Mistral) 0.59 0.7 0.67 0.65 0.67 0.84
CoQA(Llama) 0.61 0.74 0.76 0.76 0.77 0.92
CoQA(Flan-ul2) 0.61 0.76 0.78 0.78 0.79 0.87
CoQA(Mistral) 0.56 0.74 0.79 0.77 0.79 0.81
NQ(Llama) 0.65 0.75 0.75 0.73 0.74 0.85
NQ(Flan-ul2) 0.76 0.76 0.86 0.86 0.86 0.93
NQ(Mistral) 0.66 0.73 0.77 0.77 0.78 0.83
CNN (Pegasus) 0.51 0.67 0.73 0.72 0.72 0.77
CNN (BART) 0.51 0.60 0.52 0.48 0.54 0.57
XSUM (Pegasus) 0.51 0.58 0.69 0.70 0.71 0.73
XSUM (BART) 0.51 0.59 0.53 0.51 0.52 0.57
@@
Table 3: AUARCs on four Q&A and two summarization datasets (CNN, XSUM) using a total of five LLMs (Llama, Flan-ul2, Mistral, Pegasus, BART). Higher values are better. Best results bolded.
Dataset(LLM) # of Semantic Lexical EigenValue Eccentricity Degree Ours
Sets Similarity
TriviaQA(Llama) 0.77 0.8 0.8 0.8 0.8 0.83
TriviaQA(Flan-ul2) 0.69 0.72 0.73 0.73 0.73 0.74
TriviaQA(Mistral) 0.55 0.63 0.64 0.64 0.64 0.64
SQuAD(Llama) 0.3 0.36 0.37 0.28 0.36 0.68
SQuAD(Flan-ul2) 0.73 0.95 0.83 0.82 0.83 0.96
SQuAD(Mistral) 0.72 0.93 0.82 0.82 0.82 0.96
CoQA(Llama) 0.56 0.67 0.67 0.67 0.67 0.71
CoQA(Flan-ul2) 0.7 0.79 0.8 0.79 0.79 0.8
CoQA(Mistral) 0.46 0.62 0.64 0.63 0.64 0.61
NQ(Llama) 0.37 0.41 0.42 0.41 0.41 0.45
NQ(Flan-ul2) 0.41 0.44 0.47 0.46 0.45 0.47
NQ(Mistral) 0.32 0.38 0.40 0.40 0.39 0.42
CNN (Pegasus) 0.45 0.51 0.53 0.43 0.52 0.74
CNN (BART) 0.21 0.22 0.21 0.21 0.21 0.34
XSUM (Pegasus) 0.16 0.17 0.19 0.17 0.17 0.27
XSUM (BART) 0.21 0.22 0.20 0.21 0.22 0.35
@@ 4 Experiments
We demonstrate the efficacy of our method on question answering and summarization tasks. For summarization, we used BART-large [20] and Pegasus-large [46] and for question answering, we used Mistral-7B-Instruct-v0.2 [11], Llama-2-13b chat version [39], and Flan-ul2 models [36]. For question answering we elicited responses from these models on four datasets, namely, CoQA [32], SQuAD [31], TriviaQA [14] and Natural Questions (NQ) [18]. CoQA and SQuAD provide the context and expect the model to respond to the question based on the context, while TriviaQA and NQ do not have a context and require the model to tap into its learnt knowledge. For our experiments, we use the validation splits for all the datasets as done previously [22]. CoQA has 7983 datapoints, TriviaQA has 9960 datapoints, SQuAD has 10,600 datapoints and NQ has 7830 datapoints. For summarization, we used CNN  Daily Mail [33] and [8] and XSUM [27] datasets. We use a subset of the validation splits of both the datasets comprising of 4000 datapoints. For detecting entailment, we use deberta-large-nli model which is specialized for NLI tasks [7].
@@
We follow previous works [22], which used $1000$ datapoints for hyperparameter tuning, to train our Logistic Regression Classifier and the rest of them were used for evaluation. As such, in Table 8 in the appendix, we show that our method is quite performant even with fewer training datapoints. For each of the prompt perturbations specified above, we use five generations for each perturbation for more robust evaluation. All results are averaged over three runs. We use zero-shot prompting for the datasets with context. For TriviaQA, Flan-ul2 and Mistral-7B-Instruct-v0.2 also worked well with zero shot prompting while Llama-2-13b chat was performant with a two-shot prompt. For NQ, we used a five shot prompt. The details about the prompts used are provided in the Appendix A. We used internally hosted models to generate the responses. Thus, we used V100s GPUs for the feature extraction step once the responses were generated. The logistic regression model was trained on an intel core CPU.
We consider methods proposed in recent works [16] and [22] which are state-of-the-art as the baselines. [16] proposed computing the number of semantic sets and lexical similarity metrics from the generated outputs. [22] use eigen value, eccentricity and degree metrics inspired from spectral clustering to estimate the uncertainty of the model. We use the open source code provided by the authors of [22] for comparing with the baselines  [^1] https://github.com/zlin7/UQ-NLG/ The results are different in some cases from those reported in their paper possibly because of different random splits and different LLMs used, since we did run the provided code..

### v3

@@ Abstract
Estimating uncertainty or confidence in the responses of a model can be significant in evaluating trust not only in the responses, but also in the model as a whole. In this paper, we explore the problem of estimating confidence for responses of large language models (LLMs) with simply black-box or query access to them. We propose a simple and extensible framework where, we engineer novel features and train a (interpretable) model (viz. logistic regression) on these features to estimate the confidence. We empirically demonstrate that our simple framework is effective in estimating confidence of Flan-ul2, Llama-13b, Mistral-7b and GPT-4 on four benchmark Q&A tasks as well as of Pegasus-large and BART-large on two benchmark summarization tasks with it surpassing baselines by even over $10\%$ (on AUROC) in some cases. Additionally, our interpretable approach provides insight into features that are predictive of confidence, leading to the interesting and useful discovery that our confidence models built for one LLM generalize zero-shot across others on a given dataset.
@@ 1 Introduction
Having black-box access to an LLM limits the strategies one could leverage to ascertain confidence, but if the proposed strategies are effective they could be widely applied. Previous approaches [18, 24, 14] predominantly exploit the variability in the outputs for a given input prompt or based on an ensemble of prompts computing different estimators. Our approach enhances this idea where we design different ways of manipulating the input prompt and based on the variability of the answers produce values for each such manipulation. We aver to these values as features. Based on these features computed for different inputs we train a model to predict if the response was correct or incorrect. The probability of each such prediction is then the confidence that we output. Since, the models we use to produce such predictions are simple (viz. logistic regression) the confidence estimates are typically well calibrated [26]. Moreover, being interpretable we can also see which features were more crucial in the estimation. This general framework and the features we engineer are shown in Figure 1. The framework is extensible, since more features or prompt perturbations can be easily added to this framework.
@@ 2 Related Work
Table 1: Below we see examples of different prompt perturbations for a prompt from the SQuAD dataset. The color blue and strike outs indicate changes to the input prompt. i) SD does not change the prompt (hence empty cell), but using a stochastic decoding scheme samples multiple responses (four example samplings shown). PP paraphrases the prompt. SP randomly reorders some of the sentences. EFA repeats certain sentences with entities in them. SR removes stopwords. SRC checks for consistency in reasonable size random splits of the LLM response (again prompt is not perturbed). The splitting of the two sentences indicates inconsistency as depicted in red. Thus, the perturbations test an LLM in complementary ways.
@@ 3.1 Elicitation of Variable LLM Behavior
Paraphrasing (PP): In this strategy we paraphrase the context in the prompt and observe how that changes the output. An example of this is shown in Table 1. For paraphrasing, we use back translation, where we convert the original prompt into another language and translate it back into English. We use machine translation models from Helsinki-NLP on huggingface and translate the text from English to French and then back to English. This new prompt then can be used to query the LLM. Changes to the output could indicate brittleness in the LLMs original response. One could also prompt an LLM to paraphrase the responses, however, in our initial experiments, we observed that when context is involved, the model does not paraphrase the entire context and parts of it were omitted.
@@
Table 2: Percentage of prompt perturbations entailed by the original prompt for the SQuAD dataset using deberta-large-nli model. This dataset also has context unlike some of the other Q&A datasets and hence, is a more challenging case of our features to maintain semantics. As can be seen our perturbations produce the intended effect of maintaining the semantics of the original prompt in most cases.
$99.81$% $99.23$% $99.66$% $99.12$%
@@ 3.2 Featurization
Semantic Set: Based on the responses of the first strategies (run multiple times) we create semantically equivalent sets for each. A semantically equivalent set consists of outputs that are semantically the same. If a response entails another response and vice-versa, then they both are grouped under the same semantic set. The number of such sets is a feature for our model. As such, more the number of sets lower the confidence estimate. For example, if from five paraphrasings we get responses excellent, great, bad, subpar and fantastic, then the number of semantic sets would be two as excellent, great and fantastic would form one semantic set, while bad and subpar would form the other.
Lexical Similarity: We compute the average lexical similarity for outputs of each of the first five strategies (run multiple times). The similarity can be measured using standard NLP metrics such as rouge, blue score etc. The higher the lexical similarity higher the estimated confidence. We use rouge score to quantify the lexical similarity. Considering the same five paraphrasings example described above we would compute the average rouge score considering pairs of the responses and use it as a feature.
SRC Minimum Value: As mentioned above, semantic inconsistency between the two parts is measured using an NLI models contradiction probability, where one part is taken as the premise and the other the hypothesis. The highest contradiction probability amongst multiple such partitions is the feature value for this strategy. In Table 1 last row, there are only two sentences so only one split would be done and since the sentences contradict each other the NLI contradiction probability would be high or consistency would be low.
@@ 3.3 Label Creation and Confidence Estimation
Table 3: AUROCs on four Q&A and two summarization datasets (CNN, XSUM) using a total of six LLMs (Llama, Flan-ul2, Mistral, GPT-4, Pegasus, BART). Higher values are better. Best results bolded.
Dataset(LLM) # of Lexical EigenValue Eccentricity Degree SE AVC Ours
SS Similarity
TriviaQA(Llama) 0.73 0.76 0.77 0.76 0.77 0.75 0.79 0.88
TriviaQA(Flan-ul2) 0.83 0.8 0.86 0.86 0.87 0.85 0.81 0.95
TriviaQA(Mistral) 0.65 0.72 0.76 0.75 0.75 0.68 0.73 0.81 ±.$\pm.003$
TriviaQA(GPT-4) 0.89 0.91 0.91 0.92 0.91 0.92 0.94 0.96±.$\pm.007$
SQuAD(Llama) 0.65 0.72 0.74 0.58 0.72 0.61 0.61 0.83 ±.$\pm.004$
SQuAD(Flan-ul2) 0.6 0.7 0.67 0.65 0.67 0.63 0.66 0.8 ±.$\pm.007$
SQuAD(Mistral) 0.59 0.7 0.67 0.65 0.67 0.62 0.64 0.84 ±.$\pm.003$
SQuAD(GPT-4) 0.79 0.82 0.84 0.79 0.83 0.81 0.86 0.91±.$\pm.004$
CoQA(Llama) 0.61 0.74 0.76 0.76 0.77 0.64 0.78 0.92
CoQA(Flan-ul2) 0.61 0.76 0.78 0.78 0.79 0.63 0.76 0.87 ±.$\pm.001$
CoQA(Mistral) 0.56 0.74 0.79 0.77 0.79 0.59 0.75 0.81 ±.$\pm.002$
CoQA(GPT-4) 0.81 0.86 0.88 0.87 0.88 0.89 0.91 0.95±.$\pm.005$
NQ(Llama) 0.65 0.75 0.75 0.73 0.74 0.68 0.74 0.85 ±.$\pm.003$
NQ(Flan-ul2) 0.76 0.76 0.86 0.86 0.86 0.81 0.84 0.93 ±.$\pm.002$
NQ(Mistral) 0.66 0.73 0.77 0.77 0.78 0.68 0.75 0.83 ±.$\pm.003$
NQ(GPT-4) 0.81 0.85 0.85 0.85 0.88 0.89 0.9 0.93±.$\pm.003$
CNN (Pegasus) 0.51 0.67 0.73 0.72 0.72 0.55 0.73 0.77
CNN (BART) 0.51 0.60 0.52 0.48 0.54 0.53 0.5 0.57
XSUM (Pegasus) 0.51 0.58 0.69 0.70 0.71 0.54 0.71 0.73
XSUM (BART) 0.51 0.59 0.53 0.51 0.52 0.52 0.53 0.57
@@
Table 4: AUARCs on four Q&A and two summarization datasets (CNN, XSUM) using a total of six LLMs (Llama, Flan-ul2, Mistral, GPT-4, Pegasus, BART). Higher values are better. Best results bolded.
Dataset(LLM) # of Lexical EigenValue Eccentricity Degree SE AVC Ours
SS Similarity
TriviaQA(Llama) 0.77 0.8 0.8 0.8 0.8 0.79 0.8 0.83 ±.$\pm.01$
TriviaQA(Flan-ul2) 0.69 0.72 0.73 0.73 0.73 0.71 0.72 0.74 ±.$\pm.002$
TriviaQA(Mistral) 0.55 0.63 0.64 0.64 0.64 0.58 0.63 0.64 ±.$\pm.006$
TriviaQA(GPT-4) 0.8 0.84 0.84 0.84 0.82 0.84 0.85 0.89±.$\pm.004$
SQuAD(Llama) 0.3 0.36 0.37 0.28 0.36 0.36 0.31 0.68 ±.$\pm.004$
SQuAD(Flan-ul2) 0.73 0.95 0.83 0.82 0.83 0.78 0.83 0.96 ±.$\pm.003$
SQuAD(Mistral) 0.72 0.93 0.82 0.82 0.82 0.76 0.83 0.96 ±.$\pm.004$
SQuAD(GPT-4) 0.7 0.72 0.72 0.63 0.66 0.69 0.71 0.83±.$\pm.006$
CoQA(Llama) 0.56 0.67 0.67 0.67 0.67 0.61 0.66 0.71 ±.$\pm.002$
CoQA(Flan-ul2) 0.7 0.79 0.8 0.79 0.79 0.73 0.77 0.8 ±.$\pm.005$
CoQA(Mistral) 0.46 0.62 0.64 0.63 0.64 0.51 0.62 0.61 ±.$\pm.003$
CoQA(GPT-4) 0.68 0.73 0.72 0.73 0.74 0.72 0.76 0.86±.$\pm.011$
NQ(Llama) 0.37 0.41 0.42 0.41 0.41 0.39 0.42 0.45 ±.$\pm.006$
NQ(Flan-ul2) 0.41 0.44 0.47 0.46 0.45 0.44 0.45 0.47 ±.$\pm.007$
NQ(Mistral) 0.32 0.38 0.40 0.40 0.39 0.36 0.39 0.42 ±.$\pm.007$
NQ(GPT-4) 0.69 0.73 0.74 0.74 0.74 0.73 0.72 0.79±.$\pm.007$
CNN (Pegasus) 0.45 0.51 0.53 0.43 0.52 0.48 0.47 0.74 ±.$\pm.004$
CNN (BART) 0.21 0.22 0.21 0.21 0.21 0.23 0.23 0.34
XSUM (Pegasus) 0.16 0.17 0.19 0.17 0.17 0.21 0.19 0.27
XSUM (BART) 0.21 0.22 0.20 0.21 0.22 0.23 0.22 0.35
Table 5: ECEs on four Q&A and two summarization datasets (CNN, XSUM) using a total of six LLMs (Llama, Flan-ul2, Mistral, GPT-4, Pegasus, BART). Lower values are better. Best results bolded.
Dataset(LLM) # of Lexical EigenValue Eccentricity Degree SE AVC Ours
SS Similarity
TriviaQA(Llama) 0.13 0.12 0.11 0.11 0.1 0.12 0.09 0.04
TriviaQA(Flan-ul2) 0.06 0.07 0.05 0.05 0.05 0.07 0.06 0.01
TriviaQA(Mistral) 0.17 0.12 0.1 0.1 0.11 0.16 0.11 0.05
TriviaQA(GPT-4) 0.07 0.08 0.09 0.09 0.08 0.09 0.03 0.01
SQuAD(Llama) 0.15 0.12 0.1 0.24 0.13 0.18 0.18 0.04
SQuAD(Flan-ul2) 0.17 0.09 0.13 0.14 0.14 0.17 0.16 0.06
SQuAD(Mistral) 0.2 0.12 0.14 0.15 0.14 0.17 0.15 0.04
SQuAD(GPT-4) 0.11 0.09 0.08 0.19 0.07 0.1 0.11 0.02
CoQA(Llama) 0.16 0.1 0.08 0.09 0.09 0.18 0.09 0.02
CoQA(Flan-ul2) 0.15 0.11 0.09 0.09 0.09 0.17 0.08 0.03
CoQA(Mistral) 0.18 0.1 0.07 0.09 0.07 0.21 0.09 0.05
CoQA(GPT-4) 0.11 0.09 0.08 0.08 0.08 0.06 0.05 0.02
NQ(Llama) 0.13 0.08 0.08 0.09 0.09 0.12 0.08 0.04
NQ(Flan-ul2) 0.1 0.09 0.06 0.06 0.06 0.06 0.05 0.02
NQ(Mistral) 0.15 0.09 0.11 0.1 0.09 0.12 0.09 0.05
NQ(GPT-4) 0.1 0.05 0.05 0.06 0.06 0.09 0.06 0.02
CNN (Pegasus) 0.19 0.16 0.11 0.12 0.12 0.19 0.09 0.07
CNN (BART) 0.51 0.19 0.26 0.29 0.25 0.26 0.24 0.19
XSUM (Pegasus) 0.21 0.2 0.15 0.13 0.11 0.21 0.11 0.09
XSUM (BART) 0.26 0.22 0.24 0.27 0.26 0.25 0.23 0.2
Table 6: Results with different number of decodings (for each of the features) using our method. Five decodings correspond to results in the paper. As can be seen reducing to three decodings our approach still maintains performance.
Dataset(LLM) Our AUROC Our AUROC Our AUARC Our AUARC Our ECE Our ECE
5 decodings 3 decodings 5 decodings 3 decodings 5 decodings 3 decodings
TriviaQA(Llama) 0.88 0.86 0.83 0.81 0.04 0.05
TriviaQA(Flan-ul2) 0.95 0.94 0.74 0.72 0.01 0.02
TriviaQA(Mistral) 0.81 0.81 0.64 0.65 0.05 0.05
TriviaQA(GPT-4) 0.96 0.93 0.89 0.86 0.01 0.02
SQuAD(Llama) 0.83 0.81 0.68 0.65 0.04 0.06
SQuAD(Flan-ul2) 0.8 0.8 0.96 0.94 0.06 0.08
SQuAD(Mistral) 0.84 0.82 0.96 0.93 0.04 0.05
SQuAD(GPT-4) 0.91 0.89 0.83 0.8 0.02 0.03
CoQA(Llama) 0.92 0.91 0.71 0.69 0.02 0.03
CoQA(Flan-ul2) 0.87 0.85 0.8 0.78 0.03 0.05
CoQA(Mistral) 0.81 0.8 0.61 0.6 0.05 0.06
CoQA(GPT-4) 0.95 0.92 0.86 0.84 0.02 0.03
NQ(Llama) 0.85 0.83 0.45 0.44 0.04 0.06
NQ(Flan-ul2) 0.93 0.91 0.47 0.45 0.02 0.03
NQ(Mistral) 0.83 0.81 0.42 0.4 0.05 0.06
NQ(GPT-4) 0.93 0.91 0.79 0.77 0.02 0.03
CNN (Pegasus) 0.77 0.75 0.74 0.72 0.07 0.09
CNN (BART) 0.57 0.55 0.34 0.33 0.19 0.21
XSUM (Pegasus) 0.73 0.71 0.27 0.25 0.09 0.11
XSUM (BART) 0.57 0.55 0.35 0.33 0.2 0.22
Table 7: Below we see how the AUROC, AUARC and ECE values vary with different number of samples used to train our logistic regression model for the Q&A datasets. As can be seen our uncertainty estimation procedure is performant even with fewer samples for training.
Dataset LLM $250$ samples $500$ samples $1000$ samples
TriviaQA Llama 0.83, 0.80, 0.05 0.86, 0.81, 0.04 0.88, 0.83, 0.04
Flan-ul2 0.95, 0.73, 0.03 0.95, 0.74, 0.02 0.95, 0.74, 0.01
Mistral 0.80, 0.63, 0.06 0.80, 0.63, 0.06 0.81, 0.64, 0.05
GPT-4 0.92, 0.85, 0.02 0.94, 0.88, 0.01 0.96, 0.89, 0.01
SQuAD Llama 0.8, 0.65, 0.07 0.81, 0.66, 0.05 0.83, 0.68, 0.04
Flan-ul2 0.76, 0.91, 0.07 0.78, 0.94, 0.06 0.8, 0.96, 0.06
Mistral 0.79, 0.90, 0.06 0.81, 0.93, 0.05 0.84, 0.96, 0.04
GPT-4 0.89, 0.81, 0.04 0.9, 0.82, 0.02 0.91, 0.83, 0.02
CoQA Llama 0.91, 0.70, 0.04 0.92, 0.71, 0.03 0.92, 0.71, 0.02
Flan-ul2 0.86, 0.79, 0.05 0.87, 0.80, 0.04 0.87, 0.80, 0.03
Mistral 0.80, 0.60, 0.07 0.81, 0.61, 0.06 0.81, 0.61, 0.05
GPT-4 0.93, 0.84, 0.04 0.94, 0.86, 0.03 0.95, 0.86, 0.02
NQ Llama 0.81, 0.4, 0.06 0.82, 0.41, 0.05 0.85, 0.45, 0.04
Flan-ul2 0.86, 0.43, 0.04 0.87, 0.45, 0.02 0.93, 0.47, 0.02
Mistral 0.80, 0.37, 0.06 0.81, 0.39, 0.06 0.83, 0.42, 0.05
GPT-4 0.9, 0.77, 0.03 0.91, 0.78, 0.04 0.93, 0.79, 0.02
@@ 4 Experiments
We demonstrate the efficacy of our method on question answering and summarization tasks. For summarization, we used BART-large [22] and Pegasus-large [49] and for question answering, we used GPT-4 [30], Mistral-7B-Instruct-v0.2 [13], Llama-2-13b chat version [42], and Flan-ul2 models [39]. For question answering we elicited responses from these models on four datasets, namely, CoQA [35], SQuAD [34], TriviaQA [16] and Natural Questions (NQ) [20]. CoQA and SQuAD provide the context and expect the model to respond to the question based on the context, while TriviaQA and NQ do not have a context and require the model to tap into its learnt knowledge. For our experiments, we use the validation splits for all the datasets as done previously [24]. CoQA has 7983 datapoints, TriviaQA has 9960 datapoints, SQuAD has 10,600 datapoints and NQ has 7830 datapoints. For summarization, we used CNN  Daily Mail [36] and [10] and XSUM [29] datasets. We use a subset of the validation splits of both the datasets comprising of 4000 datapoints. For detecting entailment, we use deberta-large-nli model which is specialized for NLI tasks [9].
@@
TriviaQA(GPT-4) SD lexical SD semantic PP lexical -
similarity set similarity
@@
CoQA(GPT-4) SD lexical EFA semantic SD semantic SP lexical
similarity set set similarity
@@
similarity set similarity similarity
NQ(GPT-4) PP lexical SD semantic SD lexical SP lexical
@@
We follow previous works [24], which used $1000$ datapoints for hyperparameter tuning, to train our Logistic Regression Classifier and the rest of them were used for evaluation. As such, in Table 7, we show that our method is quite performant even with fewer training datapoints. This observation is also consistent when our method uses the same number of queries as our baselines as reported in 11, 12. For each of the prompt perturbations specified above, we use five generations for each perturbation for more robust evaluation. All results are averaged over five runs and we report standard deviations rounded to three decimal places for our method. We use zero-shot prompting for the datasets with context. For TriviaQA, Flan-ul2, Mistral-7B-Instruct-v0.2 and GPT-4 also worked well with zero shot prompting while Llama-2-13b chat was performant with a two-shot prompt. For NQ, we used a five shot prompt. The details about the prompts used are provided in the Appendix A. We used internally hosted models to generate the responses. Thus, we used V100s GPUs for the feature extraction step once the responses were generated. The logistic regression model was trained on an intel core CPU.
We consider methods proposed in recent works [18, 24, 44] which are state-of-the-art as the baselines. [18] proposed computing the number of semantic sets, semantic entropy and lexical similarity metrics from the generated outputs. [24] use eigen value, eccentricity and degree metrics inspired from spectral clustering to estimate the uncertainty of the model. While [44] used aggregated verbalized confidence scores. We use average verbalized confidence (AVC) as that performed the best in the previous work. To be consistent with our method we average over five estimates. We use the open source code provided by the authors of [24] for comparing with the baselines  [^1] https://github.com/zlin7/UQ-NLG/ The results are different in some cases from those reported in their paper possibly because of different random splits and different LLMs used, since we did run the provided code..
Dataset Source LLM Self Target LLM 1 Target LLM 2 Target LLM 3
TriviaQA Llama 0.88 0.94 (Flan-ul2) 0.80 (Mistral) 0.94 (GPT-4)
Flan-ul2 0.94 0.87 (Llama) 0.80 (Mistral) 0.95 (GPT-4)
Mistral 0.81 0.84 (Llama) 0.91 (Flan-ul2) 0.95 (GPT-4)
GPT-4 0.96 0.85 (Llama) 0.92 (Flan-ul2) 0.80 (Mistral)
SQuAD Llama 0.83 0.81 (Flan-ul2) 0.80 (Mistral) 0.9 (GPT-4)
Flan-ul2 0.8 0.79 (Llama) 0.78 (Mistral) 0.89 (GPT-4)
Mistral 0.84 0.82 (Llama) 0.83 (Flan-ul2) 0.91 (GPT-4)
GPT-4 0.91 0.81 (Llama) 0.8 (Flan-ul2) 0.81 (Mistral)
CoQA Llama 0.92 0.79 (Flan-ul2) 0.78 (Mistral) 0.92 (GPT-4)
Flan-ul2 0.87 0.87 (Llama) 0.81 (Mistral) 0.92 (GPT-4)
Mistral 0.81 0.88 (Llama) 0.86 (Flan-ul2) 0.93 (GPT-4)
GPT-4 0.95 0.9 (Llama) 0.86 (Flan-ul2) 0.8 (Mistral)
NQ Llama 0.85 0.91 (Flan-ul2) 0.83 (Mistral) 0.91 (GPT-4)
Flan-ul2 0.93 0.83 (Llama) 0.82 (Mistral) 0.92 (GPT-4)
Mistral 0.83 0.85 (Llama) 0.90 (Flan-ul2) 0.91 (GPT-4)
GPT-4 0.93 0.84 (Llama) 0.91 (Flan-ul2) 0.81 (Mistral)
CNN Pegasus 0.77 0.57 (BART) - -
BART 0.57 0.77 (Pegasus) - -
XSUM Pegasus 0.73 0.58 (BART) - -
BART 0.57 0.71 (Pegasus) - -